# Confluent functions, Laguerre polynomials and their (generalized) bilinear integrals

### v1

@@ 2.6 Degenerate case
$\big{(}\tfrac{1-m+\theta}{2}\big{)}_{m}F_{\theta,m}(z)=z^{-m}F_{\theta,-m}(z).$ (2.30)
@@
$\frac{1}{2\pi{\rm i}}\int\limits_{[(z,0)^{+}]}{\rm e}^{t}\Big{(}1-\frac{z}{t}\Big{)}^{-a}t^{-m-1}{\rm d}t$ $=$ ${\bf F}_{-1+2a-m,m}(z),$ (2.31)
$\frac{1}{2\pi{\rm i}}\int\limits_{[(0,1)^{+}]}\exp\Big{(}{\frac{z}{t}}\Big{)}(1-t)^{-a}t^{-m-1}{\rm d}t$ $=$ $z^{-m}{\bf F}_{-1+2a+m,-m}(z).$ (2.32)
@@
${\rm e}^{t}\big{(}1-\tfrac{z}{t}\big{)}^{-a}$ $=$ $\sum_{m\in{\mathbb{Z}}}t^{m}{\bf F}_{-1+2a-m,m}(z),$ (2.33)
$\exp\big{(}{\tfrac{z}{t}}\big{)}(1-t)^{-a}$ $=$ $\sum_{m\in{\mathbb{Z}}}t^{m}{z^{-m}{\bf F}_{-1+2a+m,-m}(z)}.$ (2.34)
@@
$U_{\theta,m}(z)=$ $\frac{(-1)^{m+1}}{\Gamma(\frac{1+\theta-m}{2})}\Bigg{(}\sum_{k=1}^{m}(-1)^{k-1}\frac{(k-1)!\big{(}\frac{1+m+\theta}{2}\big{)}_{-k}}{(m-k)!}z^{-k}$ (2.35)
@@
2.7 Relationship to the modified Bessel equation
@@
=\ = $\partial_{r}^{2}+\frac{1}{r}\partial_{r}-1-\frac{\nu^{2}}{r^{2}},\qquad z=2r,\quad\alpha=2\nu.$ (2.36)
@@
$I_{\nu}(r)$ $=\frac{1}{\Gamma(\nu+1)}\Big{(}\frac{r}{2}\Big{)}^{\nu}{\rm e}^{-r}{}_{1}F_{1}\Big{(}\nu+{\frac{1}{2}};2\nu+1;2r\Big{)}$ (2.37)
@@
$K_{\nu}(r)$ $=\sqrt{\frac{\pi}{2r}}{\rm e}^{-r}{}_{2}F_{0}\Big{(}\frac{1}{2}+\nu,\frac{1}{2}-\nu;-;-\frac{1}{2r}\Big{)}$ (2.38)
@@
2.8 Bilinear integrals
Let us start with the usual bilinear integrals of Tricomi’s functions. The following identities are known, and can be found e.g. in [6] (where instead of Kummer’s and Tricomi’s confluent functions the equivalent Whittaker functions of the first and second kind are used).
Theorem 2.1.
For $|\mathrm{Re}(\alpha)|<1$, the following identities hold:
$\quad\int_{0}^{\infty}U_{\theta_{1},\alpha}(z)U_{\theta_{2},\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z$ (2.39)
@@ 2.8 Bilinear integrals
$\int_{0}^{\infty}U_{\theta,\alpha}(z)^{2}{\rm e}^{-z}z^{\alpha}{\rm d}z$ $=\frac{\pi}{\sin\pi\alpha}\Bigg{(}\frac{\psi(\frac{1+\theta+\alpha}{2})-\psi(\frac{1+\theta-\alpha}{2})}{\Gamma(\frac{1+\theta+\alpha}{2})\Gamma(\frac{1+\theta-\alpha}{2})}\Bigg{)}.$ (2.40)
@@
$\quad\int_{0}^{\infty}U_{\theta_{1},0}(z)U_{\theta_{2},0}(z){\rm e}^{-z}{\rm d}z=\frac{2}{\theta_{1}-\theta_{2}}\frac{\psi\big{(}\tfrac{1+\theta_{1}}{2}\big{)}-\psi\big{(}\tfrac{1+\theta_{2}}{2}\big{)}}{\Gamma(\frac{1+\theta_{1}}{2})\Gamma(\frac{1+\theta_{2}}{2})},\quad\theta_{1}\neq\theta_{2},$ (2.41)
@@
$\int_{0}^{\infty}U_{\theta,0}(z)^{2}{\rm e}^{-z}{\rm d}z$ $=\frac{\psi^{\prime}\big{(}\tfrac{1+\theta}{2}\big{)}}{\Gamma\big{(}\frac{1+\theta}{2}\big{)}^{2}}.$ (2.42)
@@
By (2.19) and (2.16), we have
$\quad\frac{\theta_{1}-\theta_{2}}{2}\int_{0}^{\infty}U_{\theta_{1},\alpha}(z)U_{\theta_{2},\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z$ (2.43)
@@
where the left-right derivative is defined by $f\overset{\leftrightarrow}{\partial}g\coloneqq f\partial g-(\partial f)g$. Then, using $|\mathrm{Re}(\alpha)|<1$, the connection formula (2.28) and $\tfrac{\pi z}{\sin\pi z}=\Gamma(1+z)\Gamma(1-z)$, we obtain (2.39). The formulas (2.40), (2.41) and (2.42) are obtained by applying the rule of de l’Hôpital. $\Box$
The formulas (2.39) and (2.40) remain true for $\alpha\in{\mathbb{C}}\setminus{\mathbb{Z}}$ if the integral is replaced by the generalized integral. For $\alpha\in{\mathbb{Z}}$, the generalized integral is anomalous. It can be computed via dimensional regulatization:
Theorem 2.2.
@@
$\quad\;\mathrm{gen}\int_{0}^{\infty}U_{\theta_{1},\alpha}(z)U_{\theta_{2},\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z$ (2.44)
@@
×$\times\Big{(}-\psi(|\alpha|-k)-\psi(k+1)+\tfrac{1}{2}H_{k}\big{(}\tfrac{1+\theta_{1}-|\alpha|}{2}\big{)}-\tfrac{1}{2}H_{|\alpha|-1-k}\big{(}\tfrac{3+\theta_{2}-|\alpha|}{2}+k\big{)}\Big{)};$
$\quad\mathrm{gen}\int_{0}^{\infty}U_{\theta,\alpha}(z)^{2}{\rm e}^{-z}z^{\alpha}{\rm d}z$ (2.45)
@@
The generalized integral (2.44) can be computed by dimensional regularization as outlined in Appendix A. By (2.26), the integrand is symmetric under $\alpha\leftrightarrow-\alpha$. Hence, it is sufficient to determine determine the generalized integral for $-\alpha\in{\mathbb{N}}$.
@@
$f(\alpha,z)\coloneqq\Gamma\big{(}\tfrac{1+\theta_{1}-\alpha}{2}\big{)}\Gamma\big{(}\tfrac{1+\theta_{2}-\alpha}{2}\big{)}U_{\theta_{1},\alpha}(z)U_{\theta_{2},\alpha}(z){\rm e}^{-z}z^{\alpha}.$ (2.46)
By (2.39),
$\int_{0}^{\infty}f(\alpha,z){\rm d}z=\frac{2\pi}{(\theta_{1}-\theta_{2})\sin\pi\alpha}\Bigg{(}\frac{\Gamma(\frac{1+\theta_{2}-\alpha}{2})}{\Gamma(\frac{1+\theta_{2}+\alpha}{2})}-\frac{\Gamma(\frac{1+\theta_{1}-\alpha}{2})}{\Gamma(\frac{1+\theta_{1}+\alpha}{2})}\Bigg{)}.$ (2.47)
For $m\in{\mathbb{N}}$, one then finds
$\mathrm{fp}\int_{0}^{\infty}f(\alpha,z){\mathrm{d}}z\Big{|}_{\alpha=-m}$ (2.48)
$=$ $\frac{(-1)^{m}}{\theta_{1}-\theta_{2}}\Bigg{(}\big{(}\tfrac{1+\theta_{1}-m}{2}\big{)}_{m}\Big{(}\psi\big{(}\tfrac{1+\theta_{1}+m}{2}\big{)}+\psi\big{(}\tfrac{1+\theta_{1}-m}{2}\big{)}\Big{)}-(\theta_{1}\leftrightarrow\theta_{2})\Bigg{)}.$ (2.49)
@@
$f(\alpha,z)\coloneqq z^{\alpha}\Big{(}\Gamma\big{(}\tfrac{1+\theta_{1}-\alpha}{2}\big{)}U_{\theta_{1},\alpha}(z)\Big{)}\Big{(}\Gamma\big{(}\tfrac{1+\theta_{2}-\alpha}{2}\big{)}U_{\theta_{2},\alpha}(z){\rm e}^{-z}\Big{)}$ (2.50)
Recalling that $\alpha<0$ one sees that negative powers in (2.50) come from one source: the two terms in the brackets can be rewritten using the connection formulas (2.28) and (2.29). This yields a sum of four terms, but three of them have at least a power $z^{-\alpha}$. The only term which contains a singularity at $z=0$ is the product of
$-\frac{\pi}{\sin\pi\alpha}{\bf F}_{\theta_{1},\alpha}(z)=\sum_{k=0}^{\infty}\frac{(-1)^{k}\big{(}\frac{1+\theta_{1}-\alpha}{2}\big{)}_{k}z^{k}\Gamma(-\alpha-k)}{k!},$ (2.51)
@@
$-\frac{\pi}{\sin\pi\alpha}{\bf F}_{-\theta_{2},\alpha}(-z)=\sum_{k=0}^{\infty}\frac{\big{(}\frac{1-\theta_{2}-\alpha}{2}\big{)}_{k}z^{k}\Gamma(-\alpha-k)}{k!}.$ (2.52)
@@
$f_{p}(\alpha)$ $=\sum_{k=0}^{p}\frac{(-1)^{k}\big{(}\tfrac{1+\theta_{1}+\alpha}{2}\big{)}_{k}\big{(}\tfrac{1-\theta_{2}+\alpha}{2}\big{)}_{p-k}\Gamma(-\alpha-k)\Gamma(-\alpha-p+k)}{k!\,(p-k)!}$ (2.53)
$=\sum_{k=0}^{p}\frac{(-1)^{p}\big{(}\tfrac{1+\theta_{1}+\alpha}{2}\big{)}_{k}\big{(}\tfrac{1+\theta_{2}-\alpha}{2}-p+k\big{)}_{p-k}\Gamma(-\alpha-k)\Gamma(-\alpha-p+k)}{k!\,(p-k)!},$ (2.54)
@@
$\mathrm{gen}\int_{0}^{\infty}f(-m,z)=\mathrm{fp}\int_{0}^{\infty}f(\alpha,z){\rm d}z\Big{|}_{\alpha=-m}-\partial_{\alpha}f_{m-1}(\alpha)\Big{|}_{\alpha=-m}.$ (2.55)
We next prove (2.45). Applying the rule of de l’Hôpital to the first line of the right hand side of (2.44) we obtain
$\frac{(-1)^{\alpha}}{2\Gamma\big{(}\tfrac{1+\theta+\alpha}{2}\big{)}\Gamma\big{(}\tfrac{1+\theta-\alpha}{2}\big{)}}\Bigg{(}\psi\big{(}\tfrac{1+\theta+|\alpha|}{2}\big{)}^{2}-\psi\big{(}\tfrac{1+\theta-|\alpha|}{2}\big{)}^{2}+\psi^{\prime}\big{(}\tfrac{1+\theta+\alpha}{2}\big{)}+\psi^{\prime}\big{(}\tfrac{1+\theta-\alpha}{2}\big{)}\Bigg{)}.$ (2.56)
@@
$\big{(}\tfrac{1+\theta-|\alpha|}{2}\big{)}_{k}\big{(}\tfrac{3+\theta-|\alpha|}{2}+k\big{)}_{|\alpha|-1-k}=\frac{\big{(}\tfrac{1+\theta-|\alpha|}{2}\big{)}_{|\alpha|}}{\frac{1+\theta-|\alpha|}{2}+k}$ (2.57)
to transform the last two lines of the right hand side of (2.44) into
$\frac{(-1)^{\alpha}}{\Gamma\big{(}\tfrac{1+\theta+\alpha}{2}\big{)}\Gamma\big{(}\tfrac{1+\theta-\alpha}{2}\big{)}}\sum_{k=0}^{|\alpha|-1}\frac{-\psi(|\alpha|-k)-\psi(k+1)+\frac{1}{2}H_{k}\big{(}\tfrac{1+\theta-|\alpha|}{2}\big{)}-\frac{1}{2}H_{|\alpha|-1-k}\big{(}\tfrac{3+\theta-|\alpha|}{2}+k\big{)}}{\frac{1+\theta-|\alpha|}{2}+k}.$
@@
$\sum_{k=0}^{m-1}\frac{H_{k}(z)}{z+k}=\frac{1}{2}\big{(}H_{m}^{\prime}(z)+H_{m}(z)^{2}\big{)}.$ (2.58)
Namely, using (2.58) we obtain
$\sum_{k=0}^{|\alpha|-1}\frac{H_{k}\big{(}\tfrac{1+\theta-|\alpha|}{2}\big{)}-H_{|\alpha|-1-k}\big{(}\tfrac{3+\theta-|\alpha|}{2}+k\big{)}}{\frac{1+\theta-|\alpha|}{2}+k}$ (2.59)
@@
$\Box$
Note that for special choices of the parameters $\theta_{1}$, $\theta_{2}$, $\theta$, the right-hand sides of (2.39), (2.40), (2.41), (2.42), (2.44) and (2.45) may contain terms of the form 00 divided by 00. These special choices of parameters include the case when the Tricomi functions reduce to (multiples of) Laguerre polynomials. We therefore treat this case separately in the following section. We will show that the singularities are merely apparent and that the right-hand sides of the mentioned equations also make sense for the described special choice of parameters.
3 Laguerre polynomials
3.1 Basic properties
The Laguerre polynomials are traditionally defined by a Rodriguez-type formula:
$L_{n}^{\alpha}(z):$ $=\frac{1}{n!}{\rm e}^{z}z^{-\alpha}\partial_{z}^{n}{\rm e}^{-z}z^{n+\alpha}$ (3.1)
$=(-1)^{n}\sum_{k=0}^{n}\frac{(-\alpha-n)_{n-k}z^{k}}{k!(n-k)!}.$ (3.2)
which is meaningful for all complex $\alpha\in{\mathbb{C}}$. For real $\alpha>-1$, the Laguerre polynomials are one of the sets of classical orthogonal polynomials. Their natural interval is [0,[[0$\infty$[[ 0 , ∞ [ and the weight is $w(z)={\rm e}^{-z}z^{\alpha}$.

### v2

@@ 2.6. Degenerate case
$\big(\tfrac{1-m+\theta}{2}\big)_{m}F_{\theta,m}(z)=z^{-m}F_{\theta,-m}(z).$ (38)
@@
$\frac{1}{2\pi{\rm i}}\int\limits_{[(z,0)^{+}]}{\rm e}^{t}\Big(1-\frac{z}{t}\Big)^{-a}t^{-m-1}{\rm d}t$ $=$ ${\bf F}_{-1+2a-m,m}(z),$ (39)
$\frac{1}{2\pi{\rm i}}\int\limits_{[(0,1)^{+}]}\exp\Big({\frac{z}{t}}\Big)(1-t)^{-a}t^{-m-1}{\rm d}t$ $=$ $z^{-m}{\bf F}_{-1+2a+m,-m}(z).$ (40)
@@
${\rm e}^{t}\big(1-\tfrac{z}{t}\big)^{-a}$ $=$ $\sum_{m\in{\mathbb{Z}}}t^{m}{\bf F}_{-1+2a-m,m}(z),$ (41)
$\exp\big({\tfrac{z}{t}}\big)(1-t)^{-a}$ $=$ $\sum_{m\in{\mathbb{Z}}}t^{m}{z^{-m}{\bf F}_{-1+2a+m,-m}(z)}.$ (42)
@@
$U_{\theta,m}(z)=$ $\frac{(-1)^{m+1}}{\Gamma(\frac{1+\theta-m}{2})}\Bigg(\sum_{k=1}^{m}(-1)^{k-1}\frac{(k-1)!\big(\frac{1+m+\theta}{2}\big)_{-k}}{(m-k)!}z^{-k}$ (43)
@@
2.7. Relationship to the modified Bessel equation
@@
[…]
@@
$I_{\nu}(r)$ $=\frac{1}{\Gamma(\nu+1)}\Big(\frac{r}{2}\Big)^{\nu}{\rm e}^{-r}{}_{1}F_{1}\Big(\nu+{\frac{1}{2}};2\nu+1;2r\Big)$ (45)
@@
$K_{\nu}(r)$ $=\sqrt{\frac{\pi}{2r}}{\rm e}^{-r}{}_{2}F_{0}\Big(\frac{1}{2}+\nu,\frac{1}{2}-\nu;-;-\frac{1}{2r}\Big)$ (46)
@@
2.8. Bilinear integrals
Let us start with the usual bilinear integrals of Tricomi’s functions. The following identities are known, and can be found e.g. in [4] (where instead of Kummer’s and Tricomi’s confluent functions the equivalent Whittaker functions of the first and second kind are used). Since our paper contains many comparably long formulas, it is convenient to introduce the notation
$f(x,y)\pm(x\leftrightarrow y)\coloneqq f(x,y)\pm f(y,x)$ (47)
for some expression $f(x,y)$.
For $|\Re(\alpha)|<1$, the following identities hold:
$\quad\int_{0}^{\infty}U_{\theta_{1},\alpha}(z)U_{\theta_{2},\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z$ (48)
@@ 2.8. Bilinear integrals
$\int_{0}^{\infty}U_{\theta,\alpha}(z)^{2}{\rm e}^{-z}z^{\alpha}{\rm d}z$ $=\frac{\pi}{\sin\pi\alpha}\Bigg(\frac{\psi(\frac{1+\theta+\alpha}{2})-\psi(\frac{1+\theta-\alpha}{2})}{\Gamma(\frac{1+\theta+\alpha}{2})\Gamma(\frac{1+\theta-\alpha}{2})}\Bigg).$ (49)
@@
$\quad\int_{0}^{\infty}U_{\theta_{1},0}(z)U_{\theta_{2},0}(z){\rm e}^{-z}{\rm d}z=\frac{2}{\theta_{1}-\theta_{2}}\frac{\psi\big(\tfrac{1+\theta_{1}}{2}\big)-\psi\big(\tfrac{1+\theta_{2}}{2}\big)}{\Gamma(\frac{1+\theta_{1}}{2})\Gamma(\frac{1+\theta_{2}}{2})},\quad\theta_{1}\neq\theta_{2},$ (50)
@@
$\int_{0}^{\infty}U_{\theta,0}(z)^{2}{\rm e}^{-z}{\rm d}z$ $=\frac{\psi^{\prime}\big(\tfrac{1+\theta}{2}\big)}{\Gamma\big(\frac{1+\theta}{2}\big)^{2}}.$ (51)
@@
By (27) and (24), we have
$\quad\frac{\theta_{1}-\theta_{2}}{2}\int_{0}^{\infty}U_{\theta_{1},\alpha}(z)U_{\theta_{2},\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z$ (52)
@@
where the left-right derivative is defined by $f\overset{\leftrightarrow}{\partial}g\coloneqq f\partial g-(\partial f)g$. Then, using $|\Re(\alpha)|<1$, the connection formula (36) and $\tfrac{\pi z}{\sin\pi z}=\Gamma(1+z)\Gamma(1-z)$, we obtain (48). The formulas (49), (50) and (51) are obtained by applying the rule of de l’Hôpital. ∎
The formulas (48) and (49) remain true for $\alpha\in{\mathbb{C}}\setminus{\mathbb{Z}}$ if the integral is replaced by the generalized integral. For $\alpha\in{\mathbb{Z}}$, the generalized integral is anomalous. It can be computed via dimensional regulatization:
@@
$\quad\;\mathrm{gen}\int_{0}^{\infty}U_{\theta_{1},\alpha}(z)U_{\theta_{2},\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z$ (53)
@@
$\times\Big(-\psi(|\alpha|-k)-\psi(k+1)+\tfrac{1}{2}H_{k}\big(\tfrac{1+\theta_{1}-|\alpha|}{2}\big)-\tfrac{1}{2}H_{|\alpha|-1-k}\big(\tfrac{3+\theta_{2}-|\alpha|}{2}+k\big)\Big);$
$\quad\mathrm{gen}\int_{0}^{\infty}U_{\theta,\alpha}(z)^{2}{\rm e}^{-z}z^{\alpha}{\rm d}z$ (54)
@@
The generalized integral (53) can be computed by dimensional regularization as outlined in Appendix A, to which we refer the reader for details on the general scheme. By (34), the integrand is symmetric under $\alpha\leftrightarrow-\alpha$. Hence, it is sufficient to determine the generalized integral for $\alpha\in{\mathbb{N}}$.
@@
$f(\alpha,z)\coloneqq\Gamma\big(\tfrac{1+\theta_{1}-\alpha}{2}\big)\Gamma\big(\tfrac{1+\theta_{2}-\alpha}{2}\big)U_{\theta_{1},\alpha}(z)U_{\theta_{2},\alpha}(z){\rm e}^{-z}z^{\alpha}.$ (55)
By (48),
$\int_{0}^{\infty}f(\alpha,z){\rm d}z=\frac{2\pi}{(\theta_{1}-\theta_{2})\sin\pi\alpha}\Bigg(\frac{\Gamma(\frac{1+\theta_{2}-\alpha}{2})}{\Gamma(\frac{1+\theta_{2}+\alpha}{2})}-\frac{\Gamma(\frac{1+\theta_{1}-\alpha}{2})}{\Gamma(\frac{1+\theta_{1}+\alpha}{2})}\Bigg).$ (56)
For $m\in{\mathbb{N}}$, one then finds for the finite part (cf. Appendix A)
$\mathrm{fp}\int_{0}^{\infty}f(\alpha,z){\mathrm{d}}z\Big|_{\alpha=-m}$ (57)
$=$ $\frac{(-1)^{m}}{\theta_{1}-\theta_{2}}\Bigg(\big(\tfrac{1+\theta_{1}-m}{2}\big)_{m}\Big(\psi\big(\tfrac{1+\theta_{1}+m}{2}\big)+\psi\big(\tfrac{1+\theta_{1}-m}{2}\big)\Big)-(\theta_{1}\leftrightarrow\theta_{2})\Bigg).$ (58)
@@
$f(\alpha,z)\coloneqq z^{\alpha}\Big(\Gamma\big(\tfrac{1+\theta_{1}-\alpha}{2}\big)U_{\theta_{1},\alpha}(z)\Big)\Big(\Gamma\big(\tfrac{1+\theta_{2}-\alpha}{2}\big)U_{\theta_{2},\alpha}(z){\rm e}^{-z}\Big)$ (59)
Recalling that $\alpha<0$ one sees that negative powers in (59) come from one source: the two terms in the brackets can be rewritten using the connection formulas (36) and (37). This yields a sum of four terms, but three of them have at least a power $z^{-\alpha}$. The only term which contains a singularity at $z=0$ is the product of
$-\frac{\pi}{\sin\pi\alpha}{\bf F}_{\theta_{1},\alpha}(z)=\sum_{k=0}^{\infty}\frac{(-1)^{k}\big(\frac{1+\theta_{1}-\alpha}{2}\big)_{k}z^{k}\Gamma(-\alpha-k)}{k!},$ (60)
@@
$-\frac{\pi}{\sin\pi\alpha}{\bf F}_{-\theta_{2},\alpha}(-z)=\sum_{k=0}^{\infty}\frac{\big(\frac{1-\theta_{2}-\alpha}{2}\big)_{k}z^{k}\Gamma(-\alpha-k)}{k!}.$ (61)
@@
$f_{p}(\alpha)$ $=\sum_{k=0}^{p}\frac{(-1)^{k}\big(\tfrac{1+\theta_{1}+\alpha}{2}\big)_{k}\big(\tfrac{1-\theta_{2}+\alpha}{2}\big)_{p-k}\Gamma(-\alpha-k)\Gamma(-\alpha-p+k)}{k!\,(p-k)!}$ (62)
$=\sum_{k=0}^{p}\frac{(-1)^{p}\big(\tfrac{1+\theta_{1}+\alpha}{2}\big)_{k}\big(\tfrac{1+\theta_{2}-\alpha}{2}-p+k\big)_{p-k}\Gamma(-\alpha-k)\Gamma(-\alpha-p+k)}{k!\,(p-k)!},$ (63)
@@
$\mathrm{gen}\int_{0}^{\infty}f(-m,z)=\mathrm{fp}\int_{0}^{\infty}f(\alpha,z){\rm d}z\Big|_{\alpha=-m}-\partial_{\alpha}f_{m-1}(\alpha)\Big|_{\alpha=-m}.$ (64)
We next prove (54). Applying the rule of de l’Hôpital to the first line of the right hand side of (53) we obtain
$\frac{(-1)^{\alpha}}{2\Gamma\big(\tfrac{1+\theta+\alpha}{2}\big)\Gamma\big(\tfrac{1+\theta-\alpha}{2}\big)}\Bigg(\psi\big(\tfrac{1+\theta+|\alpha|}{2}\big)^{2}-\psi\big(\tfrac{1+\theta-|\alpha|}{2}\big)^{2}+\psi^{\prime}\big(\tfrac{1+\theta+\alpha}{2}\big)+\psi^{\prime}\big(\tfrac{1+\theta-\alpha}{2}\big)\Bigg).$ (65)
@@
$\big(\tfrac{1+\theta-|\alpha|}{2}\big)_{k}\big(\tfrac{3+\theta-|\alpha|}{2}+k\big)_{|\alpha|-1-k}=\frac{\big(\tfrac{1+\theta-|\alpha|}{2}\big)_{|\alpha|}}{\frac{1+\theta-|\alpha|}{2}+k}$ (66)
to transform the last two lines of the right hand side of (53) into
$\frac{(-1)^{\alpha}}{\Gamma\big(\tfrac{1+\theta+\alpha}{2}\big)\Gamma\big(\tfrac{1+\theta-\alpha}{2}\big)}$ $\sum_{k=0}^{|\alpha|-1}\frac{1}{\frac{1+\theta-|\alpha|}{2}+k}\Bigg(-\psi(|\alpha|-k)-\psi(k+1)$ (67)
$+\frac{1}{2}H_{k}\big(\tfrac{1+\theta-|\alpha|}{2}\big)-\frac{1}{2}H_{|\alpha|-1-k}\big(\tfrac{3+\theta-|\alpha|}{2}+k\big)\Bigg).$
@@
$\sum_{k=0}^{m-1}\frac{H_{k}(z)}{z+k}=\frac{1}{2}\big(H_{m}^{\prime}(z)+H_{m}(z)^{2}\big).$ (68)
Namely, using (68) we obtain
$\sum_{k=0}^{|\alpha|-1}\frac{H_{k}\big(\tfrac{1+\theta-|\alpha|}{2}\big)-H_{|\alpha|-1-k}\big(\tfrac{3+\theta-|\alpha|}{2}+k\big)}{\frac{1+\theta-|\alpha|}{2}+k}$ (69)
@@
Note that for special choices of the parameters $\theta_{1}$, $\theta_{2}$, $\theta$, the right-hand sides of (48), (49), (50), (51), (53) and (54) may contain terms of the form 0 divided by 0. These special choices of parameters include the case when the Tricomi functions reduce to (multiples of) Laguerre polynomials. We therefore treat this case separately in the following section. We will show that the singularities are merely apparent and that the right-hand sides of the mentioned equations also make sense for the described special choice of parameters.
3. Laguerre polynomials
3.1. Basic properties
We start with a concise overview of basic properties of Laguerre polynomials. They can be found in various sources such as [28, 1].
The Laguerre polynomials are traditionally defined by a Rodrigues-type formula:
$L_{n}^{\alpha}(z):$ $=\frac{1}{n!}{\rm e}^{z}z^{-\alpha}\partial_{z}^{n}{\rm e}^{-z}z^{n+\alpha}$ (70)
$=(-1)^{n}\sum_{k=0}^{n}\frac{(-\alpha-n)_{n-k}z^{k}}{k!(n-k)!}.$ (71)
which is meaningful for all complex $\alpha\in{\mathbb{C}}$. For real $\alpha>-1$, the Laguerre polynomials are one of the sets of classical orthogonal polynomials. Their natural interval is $0,\infty[$ and the weight is $w(z)={\rm e}^{-z}z^{\alpha}$.